Find the area of a triangle with the following dimensions: a = 4cm, b = 13cm, c = 15cm.
Answer by Alan3354(21583)

Use Heron's Law:
Area = sqrt(s(s-a)*(s-b)*(s-c)) where s = perimeter/2 = 16
A = sqrt(16*12*3*1) = sqrt(576)
Area = 24 sq cm